A shop is selling notebooks for £2.50 each.
They have a buy one get one half price offer on.
How many notebooks can be bought for £22?
1
Expert's answer
2019-09-17T11:37:13-0400

A pair of notebooks costs 2.50+2.50*0.5=£3.75.

5 pairs of notebooks cost 3.75*5=£18.75.

5 pairs of notebooks plus 1 notebook cost 18.75+2.50=£21.25.

So, 5*2+1=11 notebooks can be bought for £22 and the rest will be £0.75.